Here is a question I need help with, how do I create an equation to solve it?
Assume that one gallon of paint covers 300 square feet and that regular quality is $20 per gallon and supreme quality is $40 per gallon. Painter A is more reputable but charges $50 per hour. Painter B charges only $30 per hour. After how many hours would the two options break even? Which would be the better option, use regular quality paint and Painter A or supreme quality paint and Painter B?

Answer by addingup(3677)

20+50x = 40+30x
20x = 20
x = 1 at 1 hour they are both the same.
Over 1 hour B is cheaper. Let's prove it, make x=2
A: 20+50(2) = 120
B: 40+30(2) = 100 cheaper